### v1

# Latent Adaptive Planner for Dynamic Manipulation

Abstract
This paper presents Latent Adaptive Planner (LAP), a novel approach for dynamic nonprehensile manipulation tasks that formulates planning as latent space inference, effectively learned from human demonstration videos. Our method addresses key challenges in visuomotor policy learning through a principled variational replanning framework that maintains temporal consistency while efficiently adapting to environmental changes. LAP employs Bayesian updating in latent space to incrementally refine plans as new observations become available, striking an optimal balance between computational efficiency and real-time adaptability. We bridge the embodiment gap between humans and robots through model-based proportional mapping that regenerates accurate kinematic-dynamic joint states and object positions from human demonstrations. Experimental evaluations across multiple complex manipulation benchmarks demonstrate that LAP achieves state-of-the-art performance, outperforming existing approaches in success rate, trajectory smoothness, and energy efficiency, particularly in dynamic adaptation scenarios. Our approach enables robots to perform complex interactions with human-like adaptability while providing an expandable framework applicable to diverse robotic platforms using the same human demonstration videos.
Keywords: Imitation Learning, Robotics, Dynamic Nonprehensile Manipulation, Latent Space Planning, Classical Variational Bayes, Video-based Skill Acquisition
1 Introduction
Dynamic manipulation represents a sophisticated robotic capability that involves controlling objects through rapid changes in contact states, inertial forces, impacts, and complex physical interactions [1, 2]. While everyday human actions such as throwing, catching, and rapid object transfers all rely on dynamic physical interactions, current robotics research has predominantly focused on static or quasi-static manipulation [3, 4]. This static manipulation paradigm inherently limits robots to slow speeds, conservative grasping forces with high safety factors, and restricted motion ranges. Achieving true dynamic manipulation requires real-time consideration of diverse physical properties of objects (mass, friction, elasticity) [5, 6], changes in contact states, and joint torques [7, 8, 9], yet these dynamic characteristics are difficult to model analytically and vary significantly across different objects [10, 11].
The robotics community has attempted to address these complex dynamic manipulation challenges through two primary approaches. First, physics-based modeling approaches have sought to mathematically represent dynamics precisely, but perfectly modeling real-world uncertainties and diverse object properties remains practically impossible [12]. Second, machine learning-based approaches have attempted to learn system dynamics from data, but most consider only end-effector positions while overlooking critical dynamic elements such as contact information, joint torques, and grasping forces [3, 4, 13, 14]. While learning-based approaches often utilize data collected through teleoperation, this method faces significant practical limitations: high implementation costs, logistical challenges for achieving data scale and diversity, and technical difficulties in effectively capturing complete physical interaction data [15, 16]. Although advanced teleoperation systems capable of measuring and transmitting grasping forces, contact states, and joint torques exist in specialized research settings [17], they remain insufficiently deployed for large-scale data collection. These limitations present substantial barriers to implementing human-level adaptive and efficient dynamic manipulation capabilities in robots [18].
Traditional approaches often struggle with the temporal consistency and trajectory smoothness needed for natural robotic movements, particularly when transferring skills from human demonstrations to robotic systems [19]. While recent advances in offline reinforcement learning and imitation learning have shown promise, they often fail to generalize to scenarios requiring real-time adaptation. The emergence of generative modeling techniques for sequential decision-making has opened new avenues for addressing these challenges such as sequence modeling [20] and diffusion models [3]. However, these methods exhibit limitations in handling the long-term dependencies and environmental contingencies inherent in dynamic manipulation tasks.
In this work, we introduce Latent Adaptive Planner (LAP), which formulates planning as latent space inference effectively learned from human demonstration videos. Our approach addresses key challenges in visuomotor policy learning through a principled variational replanning framework that maintains temporal consistency while efficiently adapting to environmental changes. By leveraging a latent variable that functions as an abstract plan, LAP enables more coherent long-horizon planning while maintaining the flexibility needed for real-time adaptation. The system employs Bayesian updating in latent space to incrementally refine plans as new observations become available, striking an optimal balance between computational efficiency and real-time adaptability. Furthermore, by utilizing model-based proportional mapping to regenerate accurate kinematic-dynamic joint states and object positions from human demonstration videos, our approach bridges the gap between human motion and robotic execution, capturing the adaptive and energy-efficient qualities inherent in human manipulation skills. Importantly, this regeneration methodology allows us to create versatile datasets suitable for training diverse robotic platforms on the same manipulation tasks without requiring direct robot data collection.
2 Related Works
Human Demonstration-based Robot Data Generation
Those approaches that utilize human videos offer greater scalability but face embodiment gaps between humans and robots. Some methods address this by requiring hybrid datasets with both human and robot demonstrations [21, 22]. Others extract action labels from videos through object tracking [23, 24, 25, 26] or inverse modeling techniques [27, 28, 29]. Cross-embodiment techniques like RoviAug [30] but when applied to human-to-robot transfer [31], typically still require robot data.
Recent work has explored more efficient ways to leverage human demonstrations. Phantom [32] enables zero-shot transfer of policies trained solely on human demonstrations to robot embodiments through simple data editing techniques, eliminating the need for robot data collection. Similarly, DROID [33] demonstrates that collecting diverse manipulation data across multiple environments creates more robust policies capable of generalizing to new scenarios. DexMV [34] introduced a particularly relevant approach by establishing a framework for learning dexterous manipulation from human videos through demonstration translation techniques. This approach extracts 3D hand and object poses from videos and maps human hand trajectories to robot joint torques via inverse dynamics, closely aligning with our method of processing video data to generate robot demonstrations. Our specific data generation process, which involves human pose estimation, object pose extraction, and robot-specific scaling and kinematic retargeting, is detailed in Section 3.1.
Imitation Learning and Latent Space Planning
Imitation learning enables robots to acquire skills by mimicking demonstrations, transforming complex control into supervised learning problems that map observations to actions [35], though traditional approaches like Behavior Cloning struggle with distribution shift during deployment [36]. Recent approaches have reframed decision-making as sequence modeling [20, 37] to capture temporal dependencies in state-action sequences. Diffusion models offer an alternative paradigm for imitation learning, with Chi et al. [3] generating multimodal action distributions through iterative denoising processes and Janner et al. [38] leveraging classifier-guided diffusion models.
Latent variable models address temporal consistency challenges by encoding trajectory-level information that captures long-term dependencies. Yang et al. [39], Paster et al. [40] investigate Decision Transformer’s overfitting to environment contingencies and propose latent variable solutions that encode trajectory-level information. Latent Plan Transformer [41] formulates planning as latent space inference and decouples trajectory generation from return estimation, enabling more coherent long-horizon planning. Earlier works [42, 43] propose VAE-based models [44] for temporally extended policies, but lack the adaptive replanning capabilities central to our framework. Our work builds upon these foundations but introduces classical variational Bayes learning [45, 46] for precise plan inference and principled Bayesian updating in latent space, enabling efficient plan refinement as new observations become available as detailed in Section 3.2.
3 Method
The Latent Adaptive Planner (LAP) presents a novel methodology for dynamic manipulation tasks, addressing the inherent challenges in human demonstration learning and real-time environmental adaptation. Within this framework, effective data regeneration from human demonstrations is achieved, and adaptive responses to dynamic environments in real-time are enabled. The integration of latent space representation with variational inference allows for precise mapping between demonstration data and robot execution parameters while maintaining adaptability to unpredicted environmental changes.
3.1 Robot Model Based Data Regeneration from Human Demonstration Videos
A comprehensive framework is developed to regenerate robot training data from human demonstration videos, enabling the extraction of dynamic information without direct robot interaction. As illustrated in Figure 2, the approach consists of three main components: scene state estimation, object-robot proportional mapping, and kinematic-dynamic joint state reconstruction. This data regeneration pipeline systematically transforms human demonstrations into robot execution parameters while preserving essential dynamic characteristics.
Figure 1: Robot Model-Based Data Regeneration Pipeline from Human Demonstration Videos. The pipeline consists of three main stages: (1) Scene State Estimation, which detects and tracks box objects and human pose from demonstration videos; (2) Object-Robot Proportional Mapping, which scales object dimensions and positions relative to the robot base frame; and (3) Kinematic-Dynamic Joint State Reconstruction, which maps human joint positions to robot configurations, differentiates to obtain velocities and accelerations, and computes required joint torques through inverse dynamics including external forces.
Scene State Estimation
Scene state estimation is performed to extract spatial information from video frames. This process involves detecting and tracking box objects to determine their positions and dimensions within the scene coordinate system. Additionally, human pose estimation is conducted to track the demonstrator’s joint positions, providing the motion data necessary for the subsequent mapping process.
Object-Robot Proportional Mapping
To adapt the human demonstration environment to the robot’s workspace, object-robot proportional mapping is implemented. This process transforms detected objects’ positions and dimensions to the robot’s base coordinate frame:
${}^{R}\mathbf{p}_{\mathrm{obj}}=\mathbf{T}_{S}^{R}\cdot{}^{S}\mathbf{p}_{\mathrm{obj}}$ (1)
${}^{R}\mathbf{d}_{\mathrm{obj}}=s\cdot{}^{S}\mathbf{d}_{\mathrm{obj}}$ (2)
where ${}^{R}\mathbf{p}_{\mathrm{obj}}$ represents the object position expressed in robot frame $R$, ${}^{S}\mathbf{p}_{\mathrm{obj}}$ represents the object position expressed in scene frame $S$, $\mathbf{T}_{S}^{R}$ is the transformation matrix from scene frame to robot frame, ${}^{R}\mathbf{d}_{\mathrm{obj}}$ and ${}^{S}\mathbf{d}_{\mathrm{obj}}$ denote the object dimensions expressed in the robot and scene frames respectively, and $s$ is the scaling factor determined by the ratio between the robot arm length obtained from the robot model and the human arm length measured in pixels from the video.
Kinematic-Dynamic Joint State Reconstruction
The robot’s kinematic and dynamic states are reconstructed from the human demonstration. The joint positions are obtained through direct joint position mapping between human and robot:
$\mathbf{q}=f_{\mathrm{map}}(\mathbf{q}_{\mathrm{human}})$ (3)
where $\mathbf{q}$ represents the robot joint positions and $f_{\mathrm{map}}$ is the mapping function that accounts for the different kinematic structures between human and robot.
Joint velocities and accelerations are calculated by differentiating the joint positions with respect to time using the video’s frame rate:
$\dot{\mathbf{q}}=\frac{d\mathbf{q}}{dt}\approx\frac{\mathbf{q}_{t+\Delta t}-\mathbf{q}_{t}}{\Delta t}$ (4)
$\ddot{\mathbf{q}}=\frac{d^{2}\mathbf{q}}{dt^{2}}\approx\frac{\dot{\mathbf{q}}_{t+\Delta t}-\dot{\mathbf{q}}_{t}}{\Delta t}$ (5)
where $\Delta t$ is the time interval between consecutive frames.
With the complete kinematic state ($\mathbf{q}$, $\dot{\mathbf{q}}$, $\ddot{\mathbf{q}}$), inverse dynamics is employed to compute the required joint torques for task execution:
$\bm{\tau}=\mathbf{M}(\mathbf{q})\ddot{\mathbf{q}}+\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})\dot{\mathbf{q}}+\mathbf{G}(\mathbf{q})+\mathbf{J}^{T}(\mathbf{q})\mathbf{F}_{\mathrm{ext}}$ (6)
where $\mathbf{M}(\mathbf{q})$ is the inertia matrix, $\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})$ accounts for Coriolis and centrifugal effects, $\mathbf{G}(\mathbf{q})$ represents gravitational forces, $\mathbf{J}(\mathbf{q})$ is the Jacobian matrix, and $\mathbf{F}_{\mathrm{ext}}$ denotes external forces encountered during interaction with objects. The inclusion of external forces is crucial for accurately modeling tasks involving object manipulation and environmental contact.
This comprehensive data regeneration pipeline enables the transformation of human demonstrations into robot execution data with full kinematic and dynamic information, facilitating more efficient transfer of human skills to robotic systems.
3.2 Latent Adaptive Planner (LAP)
Our Latent Adaptive Planner (LAP) introduces a powerful framework for modeling and executing complex dynamic manipulation tasks based on human demonstrations.
Model
Given a trajectory ${\mathbf{x}}$ consisting of observations and actions from demonstrations, our LAP defines a joint probability distribution:
$p_{\theta}({\mathbf{x}},{\mathbf{z}})=p({\mathbf{z}})p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ (7)
where ${\mathbf{z}}\in\mathbb{R}^{d}$ is the latent vector representing the plan, $p({\mathbf{z}})$ is the prior model, and $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ is the trajectory generator. The prior model $p({\mathbf{z}})$ is defined as an isotropic Gaussian ${\mathbf{z}}\sim\mathcal{N}(0,I_{d})$.
The trajectory generator $p_{\theta}({\mathbf{x}}|{\mathbf{z}})$ is a conditional autoregressive model that produces actions based on the current state, historical context, and the latent plan:
$p_{\theta}({\mathbf{x}}|{\mathbf{z}})=\prod_{t=1}^{T}p_{\theta}({\mathbf{x}}_{(t)}|{\mathbf{x}}_{(t-K)},...,{\mathbf{x}}_{(t-1)},{\mathbf{z}})$ (8)
where ${\mathbf{x}}_{(t)}=(o_{t},a_{t})$, and $K$ is the context window length. The observation $o_{t}$ includes box position, contact state, and previous timestep joint positions, velocities, and torques ($\mathbf{q}_{t-1}$, $\dot{\mathbf{q}}_{t-1}$, $\bm{\tau}_{t-1}$). Action $a_{t}$ consists of current timestep joint positions, velocities, and torques ($\mathbf{q}_{t}$, $\dot{\mathbf{q}}_{t}$, $\bm{\tau}_{t}$). ${\mathbf{x}}_{(0)}$ is a special learnable token. $\theta$ is implemented using a causal Transformer model [47]. The latent plan vector ${\mathbf{z}}$ controls each step of action generation through cross-attention mechanisms as $p_{\theta}(a_{t}|o_{t-K:t-1},a_{t-K:t-1},{\mathbf{z}})$. The action is assumed to follow a single-mode Gaussian distribution with fixed variance, i.e. $a_{t}\sim\mathcal{N}(g_{\theta}(o_{t-K:t},a_{t-K:t-1},{\mathbf{z}}),I_{|a|})$.
Classical Variational Bayes Learning
We employ classical variational Bayes (VB) [45, 46, 48, 49] to learn LAP. Instead of learning an inference network (as in VAEs [44]), we directly optimize the latent plan vectors (local parameters in classical VB) for each trajectory using gradient descent. This iterative instance-level optimization offers greater precision in plan inference and greater flexibility that enables efficient adaptive (re)planning.
For each trajectory ${\mathbf{x}}$, we approximate the posterior distribution by $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$, where $\bm{\mu}$ is the posterior mean vector and $\bm{\sigma}^{2}$ is the variance-covariance matrix, assumed to be diagonal for computational efficiency. These are local parameters specific to each trajectory, while $\theta$ are global parameters shared by all samples. LAP can be learned by maximizing the evidence lower bound (ELBO),
$\begin{split}&\mathcal{L}(\theta,\bm{\mu},\bm{\sigma})=\mathbb{E}_{q({\mathbf{z}}|{\mathbf{x}})}[\log p_{\theta}({\mathbf{x}}|{\mathbf{z}})]-D_{\mathrm{KL}}(q({\mathbf{z}}|{\mathbf{x}})\|p({\mathbf{z}})),\end{split}$ (9)
where ${\mathbf{z}}\sim q({\mathbf{z}}|{\mathbf{x}})$ is sampled using reparameterization trick [44].
The training procedure alternates between the optimization of local and global parameters:
  1. (a)
    For each trajectory $\{{\mathbf{x}}_{i}\}_{i=1}^{N}$ in the mini-batch, optimize the corresponding local parameters $(\bm{\mu}_{i},\bm{\sigma}_{i})$ to maximize $\mathcal{L}(\theta,\bm{\mu}_{i},\bm{\sigma}_{i})$ with $T_{\mathrm{local}}$ steps of gradient descent (we set $T_{\mathrm{local}}=16$ in experiments).
  2. (b)
    Update the global parameters $\theta$ to maximize $\nicefrac{{1}}{{N}}\sum_{i=1}^{N}\mathcal{L}(\theta,\bm{\mu}_{i},\bm{\sigma}_{i})$.
Detailed algorithms can be found in Appendix E. This learning procedure also shares the similar intuition of fast-slow learning discussed in [49] with fast learning of local parameters and slower updates of global parameters. This approach allows our model to efficiently learn an adaptive policy specific to manipulation scenarios while maintaining general knowledge about dynamics and control strategies across various tasks.
Variational Replanning
We propose a novel approach termed variational replanning that implements a principled Bayesian update procedure in the latent space, realizing the concept of planning as latent space inference [41], which significantly improves the model efficiency.
With a learned LAP and initial observation $o_{1}$ during test time, we first sample ${\mathbf{z}}\sim p({\mathbf{z}}|o_{1})\propto p({\mathbf{z}})p({\mathbf{x}}_{0:1}|{\mathbf{z}})$ using $T_{\mathrm{local}}$ steps of gradient descent as the initial plan. As new observations become available, we aim to adaptively update the latent plan ${\mathbf{z}}$ within the Bayesian framework. With replanning horizon $\Delta$, new observations ${\mathbf{x}}_{t+1:t+\Delta}$, and previously inferred posterior $q({\mathbf{z}}|{\mathbf{x}}_{0:t})=\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}^{2}_{t})$, we update our belief through Bayesian updating:
$q({\mathbf{z}}|{\mathbf{x}}_{0:t+\Delta})=\mathcal{N}(\bm{\mu}_{t+\Delta},\bm{\sigma}^{2}_{t+\Delta})\propto q({\mathbf{z}}|{\mathbf{x}}_{0:t})p({\mathbf{x}}_{t+1:t+\Delta}|{\mathbf{x}}_{0:t},{\mathbf{z}})$ (10)
Specifically, we use classical VB to optimize the local parameters $\bm{\mu}_{t+\Delta}$ and $\bm{\sigma}_{t+\Delta}$ as:
$\bm{\mu}_{t+\Delta},\bm{\sigma}_{t+\Delta}=\arg\max_{\bm{\mu},\bm{\sigma}}\mathbb{E}_{q_{t+\Delta}}[\log p_{\theta}({\mathbf{x}}_{t+1:t+\Delta}|{\mathbf{x}}_{0:t},{\mathbf{z}})]-D_{\mathrm{KL}}(q_{t+\Delta}\|q_{t}),$ (11)
where we define $q_{t+\Delta}\triangleq q({\mathbf{z}}|{\mathbf{x}}_{0:t+\Delta})$ and $q_{t}\triangleq q({\mathbf{z}}|{\mathbf{x}}_{0:t})$ for notational simplicity. This amortized computation enables efficient replanning with just $T_{\mathrm{local}}=1$ gradient step rather than the $16$ steps required for initial planning.
The essence of our approach is treating the previously inferred distribution $q({\mathbf{z}}|{\mathbf{x}}_{0:t})$ as the prior belief and updating it with a small number of gradient steps to obtain $q({\mathbf{z}}|{\mathbf{x}}_{0:t+\Delta})$. As $t+\Delta\rightarrow T$ approaches the full trajectory length, the variational posterior $\mathcal{N}(\bm{\mu}_{t+\Delta},\bm{\sigma}^{2}_{t+\Delta})$ converges to the global posterior $\mathcal{N}(\bm{\mu}_{T},\bm{\sigma}^{2}_{T})$ that would be inferred given the complete trajectory. This establishes theoretical consistency between our incremental replanning approach and the optimal plan that would be determined with complete information. Detailed algorithms can be found in Appendix E. The integrated pipeline can be found in Figure 2.
Discussion
LAP with variational replanning offers a principled middle ground between traditional planning paradigms. Open-loop planning (${\mathbf{z}}\sim p({\mathbf{z}}|o_{1})$ inferred once) provides computational efficiency but suffers from error accumulation over time. Closed-loop planning (resampling ${\mathbf{z}}\sim p({\mathbf{z}}|{\mathbf{x}}_{0:t})$ at each step) allows adaptation but at a high computational cost. Our approach incrementally updates the latent plan through Bayesian inference, treating previous distributions as prior beliefs. This maintains the adaptive benefits of closed-loop planning while preserving computational efficiency, essentially performing online learning of the latent plan as new observations arrive.
4 Experimental Results
We evaluate our LAP on a nonprehensile dynamic manipulation task of box receiving, comparing its performance against several baseline methods including rule-based planning, behavior cloning, and Diffusion Policy approaches. This task represents a challenging example of nonprehensile dynamic manipulation, where the robot must initially interact with and control objects without stable grasping, using the dynamics of the arm movement to stabilize and then secure the incoming object.
4.1 Experimental Setup
Figure 2: System Architecture for LAP Framework. The diagram illustrates our perception-planning-control pipeline. Camera input undergoes segmentation for object detection, providing box states to the Latent Adaptive Planner (LAP). LAP operates on a dual-rate hierarchy, performing updating of latent plan at 30Hz while generating motion commands at 100Hz. Reference joint positions, velocities, and torques from LAP are refined by Model Predictive Control before execution at the motor level. Joint states and contact state are fed back to the LAP for replanning. This architecture ensures smooth execution of dynamic nonprehensile manipulation tasks while respecting robot dynamics and physical constraints, enabling real-time adaptation to environmental changes.
Task Description
Box receiving constitutes the primary focus of this study, a challenging nonprehensile dynamic manipulation task where the robotic system must intercept, temporarily control, and securely grasp boxes of varying sizes, weights, and shapes handed or tossed by human operators. This task requires real-time trajectory adaptation, impact anticipation, and force modulation as the robot first manages dynamics using arm surfaces before transitioning to a stable grasp.
Hardware Setup
To evaluate generalizability, experiments were conducted on two distinct robot platforms with different sizes and arm configurations. The experimental procedure involved data regeneration processes tailored to each robot’s specific characteristics, followed by training and validation phases. Both platforms were equipped with a ZED2 stereo camera that continuously tracks object position and trajectory through real-time segmentation algorithms. Human demonstrators introduced variability by throwing boxes with diverse velocities, heights, and release angles, ensuring comprehensive training and testing conditions. Detailed specifications of the robotic platforms are provided in the Appendix B.
Baselines
We evaluated LAP against several baselines: (1) Model-based planner that determines optimal end-effector positions and arm configurations based on box dimensions and positions, using the inverse dynamics for superior tracking performance; (2) Transformer-based Behavior Cloning (BC) using a 3-layer decoder transformer (8 attention heads, batch size 12, hidden size 64, learning rate 2e-4, trained for 2500 epochs); (3) Diffusion Policy implemented with a 1D UNet architecture using 100 denoising steps based on the original repository [3]. Our LAP uses identical architecture and hyper-parameters as BC with an additional cross-attention layer, where the number of ${\mathbf{z}}$ is 16, each with dimension 64, with $T_{\mathrm{local}}=16$ for plan optimization. All learning-based methods were trained on the same dataset regenerated from human demonstrations.
Figure 3: Impact-aware retreat trajectory for box receiving motion learned with the Latent Adaptive Planner (LAP). The left panel shows a human demonstration where the subject absorbs impact by yielding their arm along the trajectory of the incoming box before returning to the nominal pose. The right panel shows the robot reproducing this compliant motion using LAP. The red curves indicate the impact-aware retreat trajectory that minimizes energy consumption during dynamic interaction.
Table 1: Performance comparison of different methods on Robot A and B with various box types. Only trials where boxes were thrown within the reachable workspace of the fixed-base robots without excessive rotation were included in the success evaluation.
Robot Method Success (%) Energy (J) Success (%) Energy (J) Success (%) Energy (J)
Box A (66×16.5×14cm, 453g) Box B (61×30.5×30.5cm, 777g) Box C (67.1×38.2×23cm, 660g)
[…]
BC $26/30$ $31.39$ $28/30$ $38.50$ $26/30$ $37.40$
Diffusion $20/30$ $42.53$ $21/30$ $45.21$ $21/30$ $44.86$
[…]
Box A (66×16.5×14cm, 453g) Box D (48.3×22.9×24.8cm, 362g) Box E (49.5×12.7×23.5cm, 365g)
[…]
Diffusion $24/30$ $21.32$ $22/30$ $20.73$ $23/30$ $19.95$
[…]
4.2 Results and Analysis
Our experiments demonstrate that LAP consistently outperforms the baselines in all metrics and box types on both robot platforms. Table 1 summarizes the quantitative results for Robot A and Robot B, respectively. Figure 3 visualizes the impact-aware retreat trajectory for the the box receiving motion learned by LAP using Robot B.
Model-based Planning
The model-based approach achieved the highest success rate among all methods examined, however, this performance came at a significant energy cost due to its consistent application of high torque values for ensuring secure grasps. This conservative control strategy, while effective for task completion, proved suboptimal from an efficiency perspective. The implementation utilized an inverse-dynamic controller to enhance tracking performance and response time. The impedance control implementation offered inherent advantages in energy efficiency through its mass-spring-damper system, which naturally produces compliant motions when receiving impact forces. However, this approach presented substantial adaptability challenges, requiring manual gain tuning for variations in box weight and impact force, a significant limitation in dynamic environments.
Learning-based Methods
Behavioral Cloning (BC) and Diffusion-based methods demonstrated some capacity to capture human-like manipulation behaviors, but exhibited lower success rates and less stable action outputs compared to other approaches. While integration with Model Predictive Control (MPC) improved motion smoothness, these methods still failed to achieve optimal performance across the evaluation metrics.
LAP Performance
LAP balanced high success rates with superior energy efficiency compared to other policy-based methods. This performance can be attributed to two key factors: efficient replanning capabilities and effective encoding of contact dynamics and human-like retreat motions in the latent variable space during training. By learning to implement retreat motions through motion planning rather than relying solely on control-based implementations, LAP achieved a balanced performance profile. Though not matching the perfect success rates of the model-based approach with MPC, LAP demonstrated significantly improved energy efficiency through learned human-like motions that effectively absorbed impact forces and minimized overall energy consumption.
5 Conclusion
Our Latent Adaptive Planner (LAP) demonstrates significant advantages in nonprehensile dynamic manipulation tasks, achieving higher success rates while consuming less energy and maintaining lower peak torques compared to existing approaches. By introducing a latent plan as a temporal abstraction of the trajectory, our method achieves consistent long-horizon behavior while maintaining the flexibility required for dynamic environments.
The performance improvements observed across different box types highlight the effectiveness of our adaptive planning approach. Furthermore, our experiments with both Robot A and Robot B confirm that LAP successfully transfers to different robot configurations. By regenerating appropriate training data from the same human demonstration videos for each platform, our method effectively adapts to the distinct kinematic and dynamic properties of different robots. This cross-platform success validates the generalizability of our approach and demonstrates that the combination of data regeneration and latent planning provides a robust framework for dynamic manipulation tasks regardless of the specific robot embodiment.
Limitations
Despite the promising results demonstrated by our Latent Adaptive Planner, several limitations warrant acknowledgment and present opportunities for future research:
First, our approach employs a simplified binary contact representation that does not capture the continuous nature of physical interactions. More critically, we approximate contact locations by assuming the most significant impact occurs at the wrist position, rather than precisely estimating the actual contact points and force distributions. This simplification fundamentally limits our ability to model the complex dynamics of manipulation tasks where contact may occur across multiple points with varying intensities and directions.
Second, to reduce kinematic complexity, we restricted our human demonstration mapping to three primary degrees of freedom (shoulder, elbow, and wrist pitch), rather than utilizing the full seven degrees of freedom available in human arm movements. This simplification excludes potentially more efficient motion patterns that would be possible with base mobility and waist joint articulation in a full humanoid configuration. By constraining our system to a fixed-base manipulator paradigm, we limit the natural whole-body coordination that humans employ during dynamic manipulation tasks, where weight shifting and torso movements significantly contribute to task performance, energy efficiency, and stability during dynamic interactions.
Third, our reliance on 2D object and human pose estimation introduces inherent limitations in spatial reasoning, as it cannot fully capture depth-dependent interactions or account for occlusions that occur during dynamic movements. This dimensional reduction potentially impacts the fidelity of learned manipulation strategies, especially in tasks requiring precise depth perception.
Addressing these limitations will require advances in both modeling approaches and sensing technologies to capture more detailed contact dynamics, improved dimensional representation, and expanded degrees of freedom in robotic execution systems.
From a generative modeling perspective, the Latent Adaptive Planner relies on a relatively simple prior model (isotropic Gaussian) that may not adequately capture the complex, multi-modal nature of expert demonstration trajectories. More structured prior models [50, 51, 52, 53] could potentially provide richer representations of the latent space, enabling more nuanced planning capabilities and improved generalization to novel scenarios. Future work will explore these advanced modeling approaches to enhance the representational capacity and performance of our framework.
Acknowledgments
Y. W. was partially supported by NSF DMS-2015577, NSF DMS-2415226, and a gift fund from Amazon. We gratefully acknowledge the support of Lambda, Inc. for providing compute resources for this project.
Appendix A Model Predictive Control (MPC)
Dynamic Model Predictive Control (MPC) provides an effective framework for safely tracking joint-space references $\hat{q}_{k},\hat{\dot{q}}_{k},\hat{\tau}_{k}$ generated by learning policies. When implemented, this approach significantly mitigates the inherently noisy characteristics of learning-based control policies while maintaining high tracking performance. By incorporating the robot’s dynamics and physical constraints into the optimization problem, the controller ensures smooth and secure motions even when the underlying learning policy produces potentially suboptimal commands.
The formulated cost function penalizes deviations between desired and actual joint states while simultaneously limiting control effort, allowing the robot to operate efficiently within specified joint position, velocity, and torque limits. This balance between precise reference tracking and constraint satisfaction ultimately leads to more natural robot movements, enhancing the overall execution quality of learned behaviors while preventing excessive energy consumption and vibrational behaviors that might otherwise occur with direct policy execution.
QP Formulation
Define the robot state at time step $k$ as
$x_{k}=\begin{bmatrix}q_{k}\\ \dot{q}_{k}\end{bmatrix},$
where $q_{k}$ and $\dot{q}_{k}$ represent the joint angles and joint velocities, respectively. Let $u_{k}$ be the control input (torque). Over a prediction horizon of length $N_{p}$, the following cost function is minimized:
$\sum_{k=i}^{i+N_{p}}\Big{[}(\hat{q}_{k}-q_{k})^{T}\,Q_{q}\,(\hat{q}_{k}-q_{k})\;+\;(\hat{\dot{q}}_{k}-\dot{q}_{k})^{T}\,Q_{\dot{q}}\,(\hat{\dot{q}}_{k}-\dot{q}_{k})\;+\;u_{k}^{T}\,Q_{u}\,u_{k}\Big{]},$
subject to the robot dynamics and constraints. The linearized (or identified) system model is given by
$x_{k+1}=A_{k}\,x_{k}+B_{k}\,u_{k}+r_{k},$
where $A_{k}$, $B_{k}$, and $r_{k}$ approximate the robot’s dynamics around a nominal operating point. The following constraints are imposed to ensure feasibility and safety:
$u_{\min}$ $\leq u_{k}\leq u_{\max},$
$q_{\min}$ $\leq q_{k}\leq q_{\max},$
$\dot{q}_{\min}$ $\leq\dot{q}_{k}\leq\dot{q}_{\max},$
$\ddot{q}_{\min}$ $\leq\ddot{q}_{k}\leq\ddot{q}_{\max},$
$q_{0}$ $=q_{\text{current}},\quad\dot{q}_{0}=\dot{q}_{\text{current}},$
where $q_{\min},q_{\max},\dot{q}_{\min},\dot{q}_{\max}$, and so forth denote predefined joint and actuator limits. Collecting the state and input variables into a decision vector
$z_{k}=\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix},$
the QP can be expressed in a compact form as
$\min_{\{z_{k}\}_{k=i}^{i+N_{p}}}$ $\sum_{k=i}^{i+N_{p}}\bigl{(}z_{k}^{T}W_{k}\,z_{k}+2\,w_{k}^{T}\,z_{k}\bigr{)},$
subject to $x_{k+1}=A_{k}\,x_{k}+B_{k}\,u_{k}+r_{k},$
$u_{\min}\leq u_{k}\leq u_{\max},$
$q_{\min}\leq q_{k}\leq q_{\max},$
$\dot{q}_{\min}\leq\dot{q}_{k}\leq\dot{q}_{\max},$
$\ddot{q}_{\min}\leq\ddot{q}_{k}\leq\ddot{q}_{\max},$
$q_{0}=q_{\text{current}},\quad\dot{q}_{0}=\dot{q}_{\text{current}}.$
The matrices $W_{k}$ and vectors $w_{k}$ encapsulate the quadratic and linear terms derived from expanding the cost function and incorporating the linearized dynamics. Solving this QP at each time step allows the robot to effectively track the learning policy’s references $\hat{q}_{k},\hat{\dot{q}}_{k}$ while honoring physical constraints and preserving safe operation.
Appendix B Robot Configurations
B.1 Actuators
Table 2: Specifications of Different Actuator Models
Specification Koala BEAR Koala BEAR Muscle Panda BEAR Panda BEAR Plus
Dimensions (mm) $63.5\times 62\times 37$ $75\times 67\times 37.5$ $113\times 113\times 49.7$ $113\times 113\times 49.7$
Weight (g) 250 285 650 925
Peak Torque (15 sec) 4.2 Nm 8 Nm 16.8 Nm 33 Nm
Peak Torque (1.5 sec) 10.5 Nm 20 Nm 33.5 Nm 67 Nm
Table 3: Actuator Configuration for Robot A
Joint Actuator Model
Body Panda BEAR Plus
Shoulder Pitch Panda BEAR Plus
Shoulder Yaw Panda BEAR
Elbow Pitch Panda BEAR Plus
Wrist Pitch Koala BEAR
Wrist Roll Koala BEAR
Table 4: Actuator Configuration for Robot B
Joint Actuator Model
Body Koala BEAR Muscle
Shoulder Pitch Koala BEAR Muscle
Shoulder Yaw Koala BEAR Muscle
Elbow Pitch Koala BEAR Muscle
Wrist Pitch Koala BEAR Muscle
Wrist Roll Koala BEAR Muscle
B.2 Robot A
Figure 4: Coordinate system of Robot A
Table 5: Robot A Configuration
Joint Frame Coordination Coordinate (m) Joint Limit (rad)
Body Yaw $z$-axis rotation (0.0, 0.0, 1.1276) $[-1.57,1.57]$
Shoulder Yaw $z$-axis rotation (0.0, $\pm$0.21, 0.117) $[-2.27,2.27]$
Shoulder Pitch $y$-axis rotation (0.0, $\pm$0.15, 0.0) $[-1.57,1.57]$
Elbow Pitch $y$-axis rotation (0.0, 0.4, 0.0) $[-1.57,2.53]$
Wrist Pitch $y$-axis rotation (0.0, 0.375, 0.0) $[-2.44,2.44]$
Wrist Roll $x$-axis rotation (0.01925, 0.0, 0.0) $[-1.57,1.57]$
B.3 Robot B
Figure 5: Coordinate system of Robot B
Table 6: Robot B Configuration
Joint Frame Coordination Coordinate (m) Joint Limit (rad)
Body Yaw $z$-axis rotation (0.0, 0.0, 1.1327) $[-1.57,1.57]$
Shoulder Pitch $y$-axis rotation (0.0, $\pm$0.19475, 0.0) $[-1.57,1.57]$
Shoulder Roll $x$-axis rotation (0.0, 0.0, 0.0) $\begin{array}[]{l}[-1.57,0.25]\end{array}$
Elbow Yaw $z$-axis rotation (0.0, 0.0, 0.0) $\begin{array}[]{l}[-0.79,1.57]\end{array}$
Elbow Pitch $y$-axis rotation (0.02, 0.0, -0.27585) $[-0.3,2.9]$
Wrist Pitch $y$-axis rotation (0.23226, 0.019701, 0.0) $[-1.15,2.2]$
Appendix C 2D Human Pose Estimation
For 2D human pose estimation, we utilized the OpenMMLab library. In the detection phase, we used the RTMDet-M model (rtmdet_m_8xb32-100e_coco-obj365-person-235e8209.pth). Following detection, we utilized HRNet-W48 (hrnet_w48_coco_wholebody_384x288_dark-f5726563_20200918.pth) for precise keypoint estimation. The HRNet architecture maintains high-resolution representations throughout the network, enabling accurate localization of body joints. This specific model was trained on the COCO-WholeBody dataset with a 384×288 input resolution and incorporates the DARK pose estimation technique for sub-pixel accuracy. The combination of these models provided reliable human pose tracking that was essential for evaluating the physical interaction capabilities of our robotic systems.
Appendix D Box Pose Estimation
For real-time object detection in our system, we implemented YOLOv8. We created a custom dataset specifically tailored to our application environment, with all annotations performed using the Computer Vision Annotation Tool (CVAT). This approach enabled precise labeling of objects of interest while maintaining consistency across the training dataset. The segmentation model achieves an impressive inference rate of 100 Hz, allowing our system to process visual information at real-time speeds necessary for responsive robotic interaction. This high-frequency detection capability proved essential for tracking dynamic objects in the environment and facilitating accurate decision-making in our experimental scenarios.
Appendix E Latent Adaptive Planner Algorithms
In this section, we provide detailed algorithms for the Latent Adaptive Planner (LAP) framework. We describe both the training procedure and the variational replanning methodology used during inference.
E.1 Classical Variational Learning of LAP
Algorithm 1 outlines the training procedure for our Latent Adaptive Planner using Classical Variational Bayes. This approach optimizes both local parameters (latent plans for individual trajectories) and global parameters (shared decoder model) in an alternating fashion.
Algorithm 1 Classical Variational Bayes Learning for LAP
0:  Trajectory dataset $\mathcal{D}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}$, where ${\mathbf{x}}_{i}=\{(o_{t}^{i},a_{t}^{i})\}_{t=1}^{T}$
0:  Trained model parameters $\theta$
1:  Initialize global parameters $\theta$ randomly
2:  repeat
3:     Sample a mini-batch of trajectories $\{{\mathbf{x}}_{i}\}_{i=1}^{B}$ from $\mathcal{D}$
4:     for each ${\mathbf{x}}_{i}$ in the mini-batch do
5:        Initialize local parameters $(\bm{\mu}_{i},\bm{\sigma}_{i})$ randomly
6:        for $j=1$ to $T_{\text{local}}$ do
7:           Sample ${\mathbf{z}}_{i}\sim\mathcal{N}(\bm{\mu}_{i},\bm{\sigma}_{i}^{2})$ using reparameterization trick
8:           Compute ELBO: $\mathcal{L}(\theta,\bm{\mu}_{i},\bm{\sigma}_{i})=\mathbb{E}_{q({\mathbf{z}}|{\mathbf{x}}_{i})}[\log p_{\theta}({\mathbf{x}}_{i}|{\mathbf{z}})]-D_{\mathrm{KL}}(q({\mathbf{z}}|{\mathbf{x}}_{i})\|p({\mathbf{z}}))$
9:           Update $(\bm{\mu}_{i},\bm{\sigma}_{i})$ with gradient ascent on $\mathcal{L}(\theta,\bm{\mu}_{i},\bm{\sigma}_{i})$
10:        end for
11:     end for
12:     Sample ${\mathbf{z}}_{i}\sim\mathcal{N}(\bm{\mu}_{i},\bm{\sigma}_{i}^{2})$ for each trajectory in mini-batch
13:     Update $\theta$ with batch gradient ascent on $\frac{1}{B}\sum_{i=1}^{B}\mathcal{L}(\theta,\bm{\mu}_{i},\bm{\sigma}_{i})$
14:  until convergence
The key insight of this algorithm is the alternating optimization between local and global parameters. For each trajectory, we first optimize the latent plan distribution parameters $(\bm{\mu}_{i},\bm{\sigma}_{i})$ using $T_{\text{local}}$ steps of gradient ascent (typically 16 steps in our implementation). After optimizing all local parameters in the mini-batch, we then perform a single update of the global parameters $\theta$ based on the average ELBO across the mini-batch.
E.2 Variational Replanning Algorithm
Algorithm 2 presents our variational replanning approach, which enables efficient adaptation during test time as new observations become available. This approach uses Bayesian updating in the latent space, treating previous distributions as prior beliefs.
Algorithm 2 Variational Replanning for LAP
0:  Trained model parameters $\theta$, replanning horizon $\Delta$, initial observation $o_{1}$
0:  Robot actions $\{a_{t}\}_{t=1}^{T}$
1:  // Initial planning
2:  Initialize $\bm{\mu}_{0}$ and $\bm{\sigma}_{0}$ randomly
3:  for $j=1$ to $T_{\text{local}}$ do
4:     Sample ${\mathbf{z}}\sim\mathcal{N}(\bm{\mu}_{0},\bm{\sigma}_{0}^{2})$ using reparameterization trick
5:     Compute objective: $\mathcal{L}_{0}=\log p_{\theta}({\mathbf{x}}_{0:1}|{\mathbf{z}})-D_{\mathrm{KL}}(\mathcal{N}(\bm{\mu}_{0},\bm{\sigma}_{0}^{2})\|p({\mathbf{z}}))$
6:     Update $(\bm{\mu}_{0},\bm{\sigma}_{0})$ with gradient ascent on $\mathcal{L}_{0}$
7:  end for
8:  Sample ${\mathbf{z}}_{0}\sim\mathcal{N}(\bm{\mu}_{0},\bm{\sigma}_{0}^{2})$
9:  Generate and execute action $a_{1}\sim p_{\theta}(a_{1}|o_{1},{\mathbf{z}}_{0})$
10:  $t\leftarrow 1$
11:  // Adaptive replanning loop
12:  while task not complete do
13:     Observe $o_{t+1}$
14:     if $t\bmod\Delta=0$ then
15:        // Variational replanning
16:        Set $\bm{\mu}_{t}\leftarrow\bm{\mu}_{t-\Delta}$, $\bm{\sigma}_{t}\leftarrow\bm{\sigma}_{t-\Delta}$
17:        for $j=1$ to $T_{\text{replan}}$ do
18:           Sample ${\mathbf{z}}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2})$ using reparameterization trick
19:           Compute objective: $\mathcal{L}_{t}=\log p_{\theta}({\mathbf{x}}_{t-\Delta+1:t+1}|{\mathbf{x}}_{0:t-\Delta},{\mathbf{z}})-D_{\mathrm{KL}}(\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2})\|\mathcal{N}(\bm{\mu}_{t-\Delta},\bm{\sigma}_{t-\Delta}^{2}))$
20:           Update $(\bm{\mu}_{t},\bm{\sigma}_{t})$ with gradient ascent on $\mathcal{L}_{t}$
21:        end for
22:        Sample ${\mathbf{z}}_{t}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t}^{2})$
23:     end if
24:     Generate action $a_{t+1}\sim p_{\theta}(a_{t+1}|{\mathbf{x}}_{0:t+1},{\mathbf{z}}_{t})$
25:     Execute action $a_{t+1}$
26:     $t\leftarrow t+1$
27:  end while
The variational replanning algorithm has two key phases. First, we perform initial planning to establish our initial belief about the latent plan. Then, during task execution, we periodically update our belief through Bayesian updating every $\Delta$ timesteps.
A crucial aspect of this approach is that we treat the previously inferred distribution $\mathcal{N}(\bm{\mu}_{t-\Delta},\bm{\sigma}_{t-\Delta}^{2})$ as a prior for the current update, as shown in the KL-divergence term of the objective function. This enables efficient adaptation while maintaining temporal consistency in the latent plan.
In practice, we set $T_{\text{replan}}=1$ for computational efficiency, which is sufficient for incremental updates given that we’re starting from a previously optimized distribution. This stands in contrast to the more expensive initial planning phase that uses $T_{\text{local}}=16$ steps.
E.3 Implementation Details
For the variational inference optimization, we use AdamW optimizer with a learning rate of 1e-3 for optimizing local parameters and 2e-4 for global parameters. The latent dimension is set to 64, and we found that diagonal covariance matrices provide a good balance between expressiveness and computational efficiency.
The replanning horizon $\Delta$ is a tunable parameter that trades off computational cost against adaptability. In our experiments, we found $\Delta=10$ to provide good results, allowing the system to adapt to changing conditions while maintaining real-time performance.